OpenStudy (anonymous):

Matthew currently has enough money to buy 45 toy cars. If the cost of each car was 10 cents less, Matthew could buy 5 more cars. How much money does Matthew have to spend on toy cars?

OpenStudy (anonymous):

let (y) be the amount of money Matthew has. Let (x) be the price of each car: y = 45(x) Now if the cars cost 10cents less, than Matthew can buy 5 more: y = (45+5)(x - 0.10) y = 50(x - 0.10) y = 50x - 5 Now you have two equations, we'll plug in 50x - 5 for (y) in the first equation: y = 45x 50x - 5 = 45x 5x - 5 = 0 5x = 5 x = 1 So the price of each car is 1 dollar, meaning that Matthew has 45 dollars, since he could originally buy 45 cars